# revtex4-1Repair the float

### v1

@@ Acknowledgements.
A.H.A. thanks to the support from Luis Aguilar, Alejandro de León, Carlos Flores, and Jair García of the Laboratorio Nacional de Visualización Científica Avanzada. M.A.G.-A. acknowledges support from cátedra Marcos Moshinsky (MM), Universidad Iberoamericana for support with the SNI grant and the numerical analysis was also carried out by Numerical Integration for Cosmological Theory and Experiments in High-energy Astrophysics (Nicte Ha) cluster at IBERO University, acquired through cátedra MM support. A.H.A, V.M. and M.A.G.-A acknowledge partial support from project ANID Vinculación Internacional FOVI220144. M.L.M.M. thanks to the SNII-Mex́ico, CONAHCyT for support.

### v2

@@ Acknowledgements.
We thank anonymous referees for thoughtful remarks and suggestions. A.H.A. thanks the support from Luis Aguilar, Alejandro de León, Carlos Flores, and Jair García of the Laboratorio Nacional de Visualización Científica Avanzada. M.A.G.-A. acknowledges support from cátedra Marcos Moshinsky (MM), Universidad Iberoamericana for support with the SNI grant and the numerical analysis was also carried out by Numerical Integration for Cosmological Theory and Experiments in High-energy Astrophysics (Nicte Ha) cluster at IBERO University, acquired through cátedra MM support. A.H.A, V.M. and M.A.G.-A acknowledge partial support from project ANID Vinculación Internacional FOVI220144. M.L.M.M. thanks to the SNII-Mex́ico, CONAHCyT for support.